OpenStudy (anonymous):

Kendra is flying a kite 35 ft. in the air. the angle of depression from the kite to kendra is 32 degrees. if kendra holds the string 3 ft off the ground, whats the distance from her hands to her kite? Round your answer to the nearest tenth.

OpenStudy (anonymous):

a 37.7 b 41.3 c 60.4 d 66.0 ft.

OpenStudy (anonymous):

Ive concluded that it isnt A.

OpenStudy (dumbcow):

 sorry the 5 should be a 3 from the triangle you can say \[\sin 32 = \frac{35-3}{d}\] \[d = \frac{32}{\sin 32}\]

OpenStudy (anonymous):

I GOT 0.16 +

OpenStudy (anonymous):

0.16+0= ??

OpenStudy (anonymous):

 this is what i have